[Hallc_running] [New Logentry] RC daily update
simona at jlab.org
simona at jlab.org
Mon Oct 8 12:45:01 EDT 2018
We got beam back yesterday at around 6:30 pm. Except for about one hour interruption during owl when Ops tuned the machine beam has been fairly good at 60 muA (when not running on Al dummy). Since 6:30 pm yesterday we completed:
--> Q2=4.4 W=2.74 x=0.4 theta_Shms=12.81
--> Q2=4.4 W=2.74 x=0.4 theta_Shms=15.81
We changed kinematics this morning (momenta and angles) to take data at Q2=3. W=3.14 x=0.25 theta_Shms=12.42. The LH2 data taking is in progress.
For a status of the data taking please check [https://hallcweb.jlab.org/wiki/index.php/Kaon_LT_Status_Page].
Tomorrow at 7 am beam will be off for RF recovery.
We will go to a short controlled access for Steve L. to check on the magnets. If you need to access the hall, check with the Hall C RC first.
Problems requiring attention:
--> SHMS rotation: when we moved to 12.42 deg the spectrometer went to 12.42 indeed but then instead of stopping there it "slided" about 0.04 deg to higher angle (on the camera). Steve Lassiter has been notified.
--> HMS dipole LN2 valve JT4: it's stuck again and not regulating anymore. Steve L. will make a change to the software this afternoon to use the bottom valve instead.
We will meet today at 4:30 pm on the second floor of the counting house.
To join the Meeting:
To join via Room System:
Video Conferencing System: bjn.vc -or-18.104.22.168
Meeting ID : 861439920
To join via phone :
+1.408.740.7256 (US (San Jose))
+1.888.240.2560 (US Toll Free))
+1.408.317.9253 (US (Primary, San Jose))
(see all numbers - http://bluejeans.com/numbers)
2) Enter Conference ID : 861439920
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3606003
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running